OpenStudy (anonymous):

What is an equation of the line with slope -3 and y intercept of 5? Please help!

OpenStudy (sleepyjess):

\(\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\\\huge\cal\color{red}W\color{orange}E\color{goldenrod}L\color{yellow}C\color{lightgreen}O\color{darkgreen}M\color{turquoise}E~~\color{royalblue}T\color{purple}O~~\color{#00bfff}{Open}\color{#11c520}{Study}\\\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\)

OpenStudy (sleepyjess):

The equation of a line is y=mx+b where m is the slope and b is the y-intercept. Does that help?

OpenStudy (anonymous):

Are u real?

OpenStudy (anonymous):

It just says human calculator so I'm not sure

OpenStudy (sleepyjess):

Yes, I am real. XD

OpenStudy (anonymous):

Ok lol

OpenStudy (sleepyjess):

Does that help you?

OpenStudy (anonymous):

But so its y=-3x+5

OpenStudy (anonymous):

Sorta

OpenStudy (sleepyjess):

Yes! That is correct!

OpenStudy (anonymous):

But so is that all I put?

OpenStudy (anonymous):

Because my options are y=-3(x+5) y-5=-3x. -3x+y=5. Y=5x-3